OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (anonymous):

-3x - 8 + 6x = 7 first combine like terms 3x-8=7 subtract 8 from each side 3x=-1 divide by 3 x= -1/3